OpenStudy (anonymous):

Alisha has a $15,000 car loan with a 6 percent interest rate that is compounded annually. How much will she have paid at the end of the five-year loan term?

OpenStudy (anonymous):

Assuming she pays it at the end of 5 years, the total cost will be the compounded value C = P (1+0.06)^5 = 15000(1.06)^5 = 15000(1.34) = 20100. Interest will be tis total cost minus 15000. If she pays it monthly or yearly, more complicated formulas have to be used as the balance declines with time.